# A causal inference framework for spatial confounding

### v11

@@ Abstract
Over the past few decades, addressing “spatial confounding” has become a major topic in spatial statistics. However, the literature has provided conflicting definitions, and many proposed solutions are tied to specific analysis models and do not address the issue of confounding as it is understood in causal inference. We offer an analysis-model-agnostic definition of spatial confounding as the existence of an unmeasured causal confounder variable with a spatial structure. We present a causal inference framework for nonparametric identification of the causal effect of a continuous exposure on an outcome in the presence of spatial confounding. In particular, we identify two critical additional assumptions that allow the use of the spatial coordinates as a proxy for the unmeasured spatial confounder – the measurability of the confounder as a function of space, which is required for conditional ignorability to hold, and the presence of a non-spatial component in the exposure, required for positivity to hold. The latter is tied to earlier work on relative scales of exposure and confounder in the context of spatial confounding but has not previously been formalized as a positivity requirement. We also propose studying a causal estimand based on a “shift intervention” that requires less stringent identifying assumptions than traditional estimands. We propose “double machine learning” (DML), a procedure by which flexible models are used to regress both the exposure and outcome variables on confounders to arrive at a causal estimator with favorable robustness properties and convergence rates, and we prove that this approach is consistent and asymptotically normal under spatial dependence. As far as we are aware, this is the first approach to spatial confounding that does not rely on restrictive parametric assumptions (such as linearity, effect homogeneity, or Gaussianity) for both identification and estimation. We demonstrate the advantages of the DML approach analytically and via extensive simulation studies. We apply our methods and reasoning to a study of the effect of fine particulate matter exposure during pregnancy on birthweight in California.
Keywords: spatial statistics, causal inference, spatial confounding, double machine learning
@@ 1 Introduction
For most of its history, the field of spatial statistics has been concerned primarily with spatial process models, where the goal is prediction and interpolation of a spatially varying outcome $Y$ (Banerjee et al., 2008; Cressie and Wikle, 2011). However, there is now increasing interest in making inferences about the causal effect of an exposure $X$, possibly also spatially varying, on $Y$. Estimation of such causal effects needs to mitigate potential spatial confounding: unmeasured variables that are thought to possess a spatial structure and influence both the exposure and outcome of interest. Researchers have attempted to clarify when existing spatial models can support valid causal inferences (Paciorek, 2010; Hodges and Reich, 2010; Schnell and Papadogeorgou, 2020; Keller and Szpiro, 2020). These conclusions have often relied on a formalism of spatial confounding that is tied to the analysis model used. A few papers have drawn from causal inference theory and methods to propose adaptations to existing spatial models (Papadogeorgou et al., 2018; Schnell and Papadogeorgou, 2020; Guan et al., 2022; Osama et al., 2019; Thaden and Kneib, 2018; Dupont et al., 2020). These approaches still rely to some degree on parametric assumptions (e.g., linearity) used in the analysis model.
@@
We take the standpoint that spatial confounding should be defined in a manner agnostic to the analysis model used, and we provide a causal framework to define and study spatial confounding for geospatial (continuous or point-referenced) data. We thoroughly investigate the nonparametric causal assumptions that license identification of causal effects of an exposure on an outcome in presence of spatial confounding. In particular, we identify two central assumptions that allow the use of spatial coordinates as a proxy for the unmeasured spatial confounder – i) the confounder needs to be a measurable function of space, and ii) the exposure needs to have non-spatial variation. The former is crucial to ensure conditional ignorability given the spatial coordinates. The latter is a requirement for positivity and can be viewed as a formalization of the concept of relative spatial scales of variation for the exposure and confounder, identified previously in Paciorek (2010). We also recommend the estimation of an interpretable “shift estimand,” or the expected change in population-level outcomes if all individuals’ exposures were shifted by a constant amount. Identification of the shift estimand requires a less stringent and more realistic positivity assumption than identification of the entire dose-response curve. If the outcome is linear in the exposure, this estimand corresponds to the exposure coefficient in a linear model, but it is interpretable even when the outcome-exposure relationship is non-linear.
@@
In light of the proposed causal framework, we review existing methods that have been used to alleviate spatial confounding and highlight how many of them fail to address the problem or do so only under restrictive assumptions. We then turn to the application of robust methods that have been extensively studied in the causal inference literature but have, to our knowledge, never been used in the spatial literature. “Double machine learning” (DML) methods use flexible models for regressions of both the outcome and exposure on covariates. We prove a novel asymptotic normality result for the resulting estimator in the presence of spatial dependence. We demonstrate empirically that these methods can improve the efficiency, robustness, and interoperability of inference about causal effects in spatial data. As far as we are aware, this is the first approach to spatial confounding that does not rely on restrictive parametric assumptions.
@@ 2 Spatial and causal confounding
The term “spatial confounding” is frequently used to explain why point estimates from spatial models may be different from point estimates from naive models like ordinary least squares (OLS) that do not use any spatial information. However, this term has rarely been defined explicitly, and we have identified at least four phenomena that are sometimes referred to as “spatial confounding”: (1) Omitted confounder bias: The existence of an unmeasured confounding variable with a spatial structure and that influences both the exposure and the outcome [e.g., Schnell and Papadogeorgou (2020)]. (2) Random effect collinearity: The change in fixed effect estimates that may come about when spatially-dependent random effects added to a regression model are collinear with the covariates [e.g., Hodges and Reich (2010)]. (3) Regularization bias: The finite-sample bias of methods that use flexible regression functions like splines or Gaussian processes (GP) to control for an unknown function of space [e.g., Dupont et al. (2020)]. (4) Concurvity: The difficulty of assessing the effect of an exposure which is, or is close to, a smooth function of space, if an arbitrary smooth function of space is also included in a regression model [discussed in Ramsay et al. (2003); Paciorek (2010)].
@@ 2.1 A brief review of causal inference concepts
Causal effects are typically defined in terms of potential outcomes or counterfactuals. Let $Y$ be an outcome and $X$ an exposure of interest. A potential or counterfactual outcome $Y(x)$ is the outcome that would have been observed if, possibly contrary to fact, the exposure had been set to value $x$, holding all else constant. Typically only one potential outcome is observed for each unit; the fact that only one of two (if $X$ is binary) or infinite (if $X$ is continuous) potential outcomes can be observed is known as the “fundamental problem of causal inference” (Rubin, 1974). The challenge, then, is to make inferences about the expectation or distribution of potential outcomes despite not being able to fully observe them.
@@
We first consider identification of the causal estimand $E[Y(x)]$. If this is identified for all $x$ then the entire dose-response curve is identified and so are causal effects such as $E[Y(x)]-E[Y(x^{\prime})]$, i.e., the expected change in potential outcomes if $X$ were set to $x$ compared to $x^{\prime}$. In general $E[Y(x)]$ is not identified from the observed data, but it can be identified with the help of measured covariates $C$ under the following assumptions:
Assumption 2.1 (Consistency).
@@
Under assumptions A2.1-A2.3, $E[Y(x)]$ can be identified as follows:
$E[Y(x)]=E[E[Y(x)|C]]=E[E[Y(x)|X,C]]=E[E[Y|X=x,C]]$ (1)
@@
$E[Y(x)]$ $=E\left[E\left[Y(x)|C\right]\right]=E\left[E\left[Y(x)\frac{I_{X=x}}{P(X=x|C)}|C\right]\right]$
$=E\left[E\left[Y\frac{I_{X=x}}{P(X=x|C)}|C\right]\right]=E\left[Y\frac{I_{X=x}}{P(X=x|C)}\right]$ (2)
@@
Unmeasured confounding is present if ignorability holds conditional on $C,U$ for measured confounders $C$ and some unmeasured confounder $U$, but $C$ alone does not suffice for ignorability to hold. That is, there is unmeasured confounding by $U$ if $Y(x)\perp X|C,U$ but $Y(x)\not\perp X|C$. In such a case, conditional ignorability does not hold when conditioning only on $C$, and the estimand is not identified by $\mathcal{P}_{obs}(Y,X,C)$.
@@ 2.2 Towards a causally informed definition of spatial confounding
With this fact in mind, the presence of spatial confounding may be seen as a boon contrasted with unmeasured confounding that has no spatial structure. Spatial information may be leveraged to capture some of the variability in the confounder and, possibly, to control for some of the confounding. In this case, it is the assumption of a spatial structure that creates the bridge between the observed data and the unmeasured confounder. In Section 3 we provide formal conditions under which spatial coordinates suffice to control for unmeasured spatial confounding, but we caution here that the mere fact that the confounder is spatially varying does not suffice to guarantee that spatial (or any) methods can mitigate or control for confounding.
@@ 3 Identification in the presence of spatial confounding
Assumption 3.1 (Ignorability conditional on unmeasured confounder $U$).
@@
Assumption 3.2 (Positivity with respect to $U$ ).
@@ 3.1 Ignorability and spatial confounding
There are two ways in which an arbitrary unmeasured confounder could fail to be equal to $g(S)$ for some measurable function $g$. First, $U$ may not be a function of space alone; this is the case whenever two individuals may share a location without sharing a value of $U$. For example, if $U$ is personal income but location is measured at the household level, then $U$ is not a function of $S$ alone – though it may depend partially on $S$ – because members of the same household may have different values of personal income.
On the other hand, even if it can be claimed that $U$ is a function of $S$, it is not necessarily a measurable function of $S$. On a continuous domain, a measurable function can be thought of as a “nearly continuous” function, or a continuous function with at most a small set of discontinuities (see “Lusin’s theorem” in Pugh (2015)). Then even a variable such as household income (as opposed to personal income) does not satisfy this criterion, because the map from spatial location to household income is likely very discontinuous, as there is non-spatial variation (e.g., between the incomes of neighboring households).
@@
Now we state our main proposition characterizing proving conditions under which conditioning on location $S$ in lieu of an unmeasured spatial confounder $U$ suffices for ignorability to hold (possibly also conditional on measured covariates $C$, omitted for conciseness).
Proposition 3.1.
@@
Assumption 3.4 is similar to the conditional ignorability Assumption 3.1. If the latter is true, the former is often a reasonable assumption; it would typically hold whenever spatial location is causally precedent to exposure and outcome but would be violated if spatial location were a collider between exposure and outcome (i.e., it is influenced by both exposure and outcome) or a descendant of (i.e., influenced directly or indirectly by) such a collider. It would also be violated if spatial location were a mediator of the $X-Y$ relationship, e.g., if $X$ affected where people chose to live which in turn affected $Y$. The assumption is likely uncontroversial in many applications, though this could be complicated in longitudinal studies, especially if there is movement of subjects. (See Elwert and Winship (2014) for a discussion of collider bias.) Assumption 3.4 may be circumvented in favor of the following stronger but possibly more intuitive assumption:
Assumption 3.5.
@@
This assumption implies that spatial location only affects the joint distribution of potential outcomes and exposure through the unmeasured confounder. This assumption is reasonable if we view $U$ as the collection of all the (unmeasured) spatial causes of $X$ and/or $Y$. Together with Assumptions 3.1 and 3.3, this is sufficient for the conclusion of Proposition 3.1.
@@ 3.2 Positivity and spatial confounding
The following proposition says that if positivity holds conditional on $U$, and $S$ is only associated with $X$ through $U$, then positivity also holds conditional on $S$, i.e., Assumption 3.6 follows from Assumptions 3.2 and 3.7:
Proposition 3.2.
Under Assumptions 3.2 and 3.7, positivity holds conditional on $S$.
@@
An important implication of positivity is that the exposure needs to have some source of non-spatial variation. If the exposure is purely a function of space, then positivity (Assumption 3.6) is violated and the causal effect cannot be identified. We note that this requirement of a non-spatial variation in the exposure has also been used in establishing consistency of specific semi-parametric spatial models (Yang et al., 2015; Dupont et al., 2020).
@@ 3.3 Estimability
First, it must be the case that $g(S)$ is regular enough, in some sense, to be estimable by the chosen strategy. No method can well-approximate an arbitrarily complex function in finite samples, therefore all methods make some assumptions about the complexity of $g(S)$. For example, spline methods assume that the function is smooth.
@@
Conversely, researchers who reason directly about the functional form of $E[Y|X,S]$ may miss important interaction or other terms needed to capture $E[Y|X,g(S)]$. Suppose the unmeasured confounder $U$ interacts with $X$ in its effect on $Y$. Then an additive model for $E[Y|X,S]$ that does not include interactions between $X$ and functions of $S$ cannot hope to control for confounding by $U$. Again, this highlights the need for flexible nonparametric estimation strategies.
@@ 3.4 Shift interventions
So far we have focused on the nonparametric identification of $E[Y(x)]$, but typically the goal of causal inference is to estimate a causal effect rather than a mean potential outcome. In the simplest case, a binary exposure $X=0,1$ admits the notion of an “average treatment effect,” defined as $E[Y(1)-Y(0)]$. When $X$ is continuous, as is often the case with environmental exposures, there is no longer just one quantity which completely specifies the average effect of interventions on $X$. One might choose to estimate the entire exposure-response curve $E[Y(x)]$ as a function of $x$, but it can be convenient to have a more concise characterization of the effect of an intervention. As we will describe in Section 4, this is typically accomplished in existing approaches by assuming linear and homogeneous treatment effects so that any contrast $E[Y(x)-Y(x^{\prime})]$ is a multiple of the exposure coefficient in a linear regression. However, when the operative parametric assumptions do not hold, this approach does not necessarily result in interpretable causal effects. While estimation of the exposure-response curve using robust causal methods of the kind we advocate is feasible (e.g., Kennedy et al. (2017)), inference is not straightforward and functional estimation may be unnecessarily complicated if a low-dimensional effect summary is sufficient.
We define a “shift intervention” as a special case of the class of interventions considered in Haneuse and Rotnitzky (2013) and Díaz Muñoz and van der Laan (2012). A shift intervention is an intervention that assigns to individual $i$ with exposure $X_{i}$ the new exposure $X_{i}+\delta$ for some constant $\delta$. Thus the expected effect of the shift intervention on individual $i$ is $E[Y_{i}(X_{i}+\delta)-Y_{i}(X_{i})]$; notably, this quantity depends on the observed exposure $X_{i}$. If we are interested in population-level effects, we may define $\Delta=E[Y(X+\delta)-Y(X)]$, simply taking the population average of individual effects, where the expectation is taken over the observed data distribution of $X$. The shift intervention effect $\Delta$ is the average effect of increasing the exposure of each individual in the population by $\delta$. This estimand allows the effect of the intervention to depend on the existing distribution of exposure in the population. For example, if the exposure only affects the outcome above a certain threshold, shift interventions might have no effect if current population levels are far below the threshold, but high effects if current population levels are near or above the threshold.
@@
The shift positivity assumption is satisfied if, whenever there is positive probability of exposure level $x$ conditional on confounder value $u$, there is also positive probability of exposure level $x+\delta$ conditional on $U=u$. In the case of spatial confounding, if Assumption 3.8 holds, and Assumption 3.7 is satisfied, then shift positivity also holds for $S$; the logic is identical to that of the case of positivity. Alternatively, we can weaken Assumption 3.6 to a version of Assumption 3.8 that replaces $U$ with of $S$. Shift positivity may be a more reasonable assumption for exposures such as air pollution, where there may be enough variation in the exposure over relatively small regions to satisfy shift positivity, but not enough exposure variation over small regions to cover the entire range of possible exposure values. This also leads to a weaker version of the ignorability assumption (Haneuse and Rotnitzky, 2013) that only enforces ignorability for the specific exposure values under consideration:
Assumption 3.9 (Shift ignorability).
For $(x,u)$ in the support of $(X,U)$, $f(Y(x+\delta)|x,u)=f(Y(x+\delta)|x+\delta,u)$.

### v12

@@ Abstract
Over the past few decades, addressing “spatial confounding” has become a major topic in spatial statistics. However, the literature has provided conflicting definitions, and many proposed solutions are tied to specific analysis models and do not address the issue of confounding as it is understood in causal inference. We offer an analysis-model-agnostic definition of spatial confounding as the existence of an unmeasured causal confounder variable with a spatial structure. We present a causal inference framework for nonparametric identification of the causal effect of a continuous exposure on an outcome in the presence of spatial confounding. In particular, we identify two critical additional assumptions that allow the use of the spatial coordinates as a proxy for the unmeasured spatial confounder – the measurability of the confounder as a function of space, which is required for conditional ignorability to hold, and the presence of a non-spatial component in the exposure, required for positivity to hold. We also propose studying a causal estimand based on a “shift intervention” that requires less stringent identifying assumptions than traditional estimands. We then turn to estimation and focus on “double machine learning” (DML), a procedure in which flexible models are used to regress both the exposure and outcome variables on confounders to arrive at a causal estimator with favorable robustness properties and convergence rates. This procedure avoids restrictive assumptions, such as linearity and effect homogeneity, which are typically made in spatial models and which can lead to bias when violated. We demonstrate the advantages of the DML approach analytically and via extensive simulation studies. We apply our methods and reasoning to a study of the effect of fine particulate matter exposure during pregnancy on birthweight in California.
@@ 1 Introduction
For most of its history, the field of geostatistics has been concerned primarily with spatial process models, where the goal is prediction and interpolation of a spatially varying outcome $Y$ (Banerjee et al., 2008; Cressie and Wikle, 2011). However, there is now increasing interest in making inferences about the causal effect of an exposure $X$, possibly also spatially varying, on $Y$. Estimation of such causal effects needs to mitigate potential spatial confounding: unmeasured variables that are thought to possess a spatial structure and influence both the exposure and outcome of interest. Researchers have attempted to clarify when existing spatial models can support valid causal inferences (Paciorek, 2010; Hodges and Reich, 2010; Schnell and Papadogeorgou, 2020; Keller and Szpiro, 2020). These conclusions have often relied on a formalism of spatial confounding that is tied to the analysis model used. A few papers have drawn from causal inference theory and methods to propose adaptations to existing spatial models (Papadogeorgou et al., 2018; Schnell and Papadogeorgou, 2020; Guan et al., 2022; Osama et al., 2019; Thaden and Kneib, 2018; Dupont et al., 2020; Bobb et al., 2022; Khan and Berrett, 2023). These approaches still rely to some degree on parametric assumptions (e.g., linearity) used in the analysis model.
@@
We take the standpoint that spatial confounding should be defined in a manner agnostic to the analysis model used, and we provide a causal framework to define and study spatial confounding for geospatial (continuous or point-referenced) data. We thoroughly investigate the nonparametric causal assumptions that license the identification of causal effects of an exposure on an outcome in the presence of spatial confounding. In particular, we identify two central assumptions that allow the use of spatial coordinates as a proxy for the unmeasured spatial confounder – i) the confounder needs to be a measurable function of space, and ii) the exposure needs to have non-spatial variation. The former is crucial to ensure conditional ignorability given the spatial coordinates. The latter is a requirement for positivity and can be viewed as a formalization of the concept of relative spatial scales of variation for the exposure and confounder, identified previously in Paciorek (2010). We also recommend the estimation of an interpretable “shift estimand,” or the expected change in population-level outcomes if all individuals’ exposures were shifted by a constant amount. Identification of the shift estimand requires a less stringent and more realistic positivity assumption than identification of the entire dose-response curve. If the outcome is linear in the exposure, this estimand corresponds to the exposure coefficient in a linear model, but it is interpretable even when the outcome-exposure relationship is non-linear.
@@
In light of the proposed causal framework, we review existing methods that have been used to alleviate spatial confounding and highlight how many of them fail to address the problem or do so only under restrictive assumptions. We then turn to the application of robust methods that have been extensively studied in the causal inference literature but have, to our knowledge, not been used in the spatial literature. “Double machine learning” (DML) methods use flexible models for regressions of both the outcome and exposure on covariates. We demonstrate empirically that these methods can improve the efficiency, robustness, and interoperability of inference about causal effects in spatial data.
@@ 2 Spatial and causal confounding
The term “spatial confounding” is frequently used to explain why point estimates from spatial models may be different from point estimates from naive models like ordinary least squares (OLS) that do not use any spatial information. However, this term has rarely been defined explicitly, and we have identified at least four phenomena that are sometimes referred to as “spatial confounding”:
  1. 1.
    Omitted confounder bias: The existence of an unmeasured confounding variable with a spatial structure and that influences both the exposure and the outcome [e.g. Schnell and Papadogeorgou (2020)].
  2. 2.
    Random effect collinearity: The change in fixed effect estimates that may come about when spatially-dependent random effects added to a regression model are collinear with the covariates [e.g. Hodges and Reich (2010)].
  3. 3.
    Regularization bias: The finite-sample bias of methods that use flexible regression functions like splines or Gaussian processes (GP) to control for an unknown function of space [e.g. Dupont et al. (2020)].
  4. 4.
    Concurvity: The difficulty of assessing the effect of an exposure which is, or is close to, a smooth function of space, if an arbitrary smooth function of space is also included in a regression model [discussed in Ramsay et al. (2003); Paciorek (2010)].
@@ 2.1 A brief review of causal inference concepts
Causal effects are typically defined in terms of potential outcomes or counterfactuals. Let $Y$ be an outcome and $X$ an exposure of interest. A potential or counterfactual outcome $Y(x)$ is the outcome that would have been observed if, possibly contrary to fact, the exposure had been set to value $x$, holding all else constant. Typically, only one potential outcome is observed for each unit; the fact that only one of two (if $X$ is binary) or infinite (if $X$ is continuous) potential outcomes can be observed is known as the “fundamental problem of causal inference” (Rubin, 1974). The challenge, then, is to make inferences about the expectation or distribution of potential outcomes despite not being able to fully observe them.
@@
We first consider the identification of the causal estimand $E[Y(x)]$. If this is identified for all $x$ then the entire dose-response curve is identified, and so are causal effects such as $E[Y(x)]-E[Y(x^{\prime})]$, i.e., the expected change in potential outcomes if $X$ were set to $x$ compared to $x^{\prime}$. In general, $E[Y(x)]$ is not identified from the observed data, but it can be identified with the help of measured covariates $C$ under the following assumptions:
A2.1 (Consistency).
@@
Under assumptions 2.1-2.3, $E[Y(x)]$ can be identified as follows:
$E[Y(x)]$ $=E[E[Y(x)|C]]=E[E[Y(x)|X,C]]$ (1)
$=E[E[Y|X=x,C]]$
@@
$E[Y(x)]$ $=E\left[E\left[Y(x)|C\right]\right]=E\left[E\left[Y(x)\frac{I_{X=x}}{P(X=x|C)}|C\right]\right]$ (2)
$=E\left[E\left[Y\frac{I_{X=x}}{P(X=x|C)}|C\right]\right]=E\left[Y\frac{I_{X=x}}{P(X=x|C)}\right]$
@@
Outcome regression and propensity score regression can be combined to generate doubly robust estimation strategies that are consistent for $E[E[Y|C,X=x]]=E\left[Y\frac{I_{X=x}}{P(X=x|C)}\right]$ if either one, but not necessarily both, of the outcome regression and propensity score models are correctly specified. We return to doubly robust estimation in Section 5.
@@ 2.2 Towards a causally informed definition of spatial confounding
With this fact in mind, the presence of spatial confounding may be seen as a boon contrasted with unmeasured confounding that has no spatial structure. Spatial information may be leveraged to capture some of the variability in the confounder and, possibly, to control for some of the confounding. In this case, it is the assumption of a spatial structure that creates the bridge between the observed data and the unmeasured confounder. In Section 3, we provide formal conditions under which spatial coordinates suffice to control for unmeasured spatial confounding, but we caution here that the mere fact that the confounder is spatially varying does not suffice to guarantee that spatial (or any) methods can mitigate or control for confounding.
@@ 3 Identification in the presence of spatial confounding
A3.1 (Ignorability conditional on unmeasured confounder $U$).
@@
A3.2 (Positivity with respect to $U$ ).
@@ 3.1 Ignorability and spatial confounding
There are two related ways in which an arbitrary unmeasured confounder could fail to be equal to $g(S)$ for some measurable function $g$. First, $U$ may not be a function of space alone; this is the case whenever two individuals may share a location without sharing a value of $U$. For example, if $U$ is personal income but location is measured at the household level, then $U$ is not a function of $S$ alone – though it may depend partially on $S$ – because members of the same household may have different values of personal income. On the other hand, even if it can be claimed that $U$ is a function of $S$, it is not necessarily a measurable function of $S$. On a continuous domain, a measurable function can be thought of as a “nearly continuous” function, or a continuous function with at most a small set of discontinuities (see “Lusin’s theorem” in Pugh (2015)). Then even a variable such as household income (as opposed to personal income) may not satisfy this criterion, if the map from spatial location to household income is highly discontinuous.
@@
Now we state our main proposition characterizing conditions under which conditioning on location $S$ in lieu of an unmeasured spatial confounder $U$ suffices for ignorability to hold (possibly also conditional on measured covariates $C$, omitted for conciseness).
Proposition 1.
@@
Assumption 3.4 is similar to the conditional ignorability Assumption 3.1. If the latter is true, the former is often a reasonable assumption; it would typically hold whenever spatial location is causally precedent to exposure and outcome but would be violated if spatial location were a collider between exposure and outcome (i.e., it is influenced by both exposure and outcome) or a descendant of (i.e., influenced directly or indirectly by) such a collider. It would also be violated if spatial location were a mediator of the $X-Y$ relationship, e.g., if $X$ affected where people chose to live, which in turn affected $Y$. The assumption is likely uncontroversial in many applications, though this could be complicated in longitudinal studies, especially if there is movement of subjects. (See Elwert and Winship (2014) for a discussion of collider bias.) Assumption 3.4 may be circumvented in favor of the following stronger but possibly more intuitive assumption:
A3.5.
@@
This assumption implies that spatial location only affects the joint distribution of potential outcomes and exposure through the unmeasured confounder. This assumption is reasonable if we view $U$ as the collection of all the (unmeasured) spatial causes of $X$ and/or $Y$. Together with Assumptions 3.1 and 3.3, this is sufficient for the conclusion of Proposition 1.
@@ 3.2 Positivity and spatial confounding
The following proposition says that if positivity holds conditional on $U=g(S)$, and $S$ is only associated with $X$ through $U$, then positivity also holds conditional on $S$, i.e., Assumption 3.6 follows from Assumptions 3.2 and 3.7:
Proposition 2.
Under Assumptions 3.2, 3.3 and 3.7, positivity holds conditional on $S$.
@@
An important implication of positivity is that the exposure needs to have some source of non-spatial variation. If the exposure is purely a function of space, then positivity (Assumption 3.6) is likely to be violated and the causal effect cannot be identified. We note that this requirement of a non-spatial variation in the exposure has also been used in establishing consistency of specific semi-parametric spatial models (Yang et al., 2015; Dupont et al., 2020; Gilbert et al., 2023).
@@ 3.3 Estimability
First, it must be the case that $g(S)$ is regular enough, in some sense, to be estimable by the chosen strategy. No method can approximate an arbitrarily complex function in finite samples, therefore all methods make some assumptions about the complexity or smoothness of $g(S)$.
@@
Conversely, researchers who reason directly about the functional form of $E[Y|X,S]$ may miss important interactions or other terms needed to capture $E[Y|X,g(S)]$. Suppose the unmeasured confounder $U$ interacts with $X$ in its effect on $Y$. Then an additive model for $E[Y|X,S]$ that does not include interactions between $X$ and functions of $S$ cannot hope to control for confounding by $U$. Again, this highlights the need for flexible nonparametric estimation strategies. We discuss these issues concretely when studying various estimators in Section 4.
@@ 3.4 Shift interventions
So far, we have focused on the nonparametric identification of $E[Y(x)]$, but typically the goal of causal inference is to estimate a causal effect rather than a mean potential outcome. In the simplest case, a binary exposure $X=0,1$ admits the notion of an “average treatment effect”, defined as $E[Y(1)-Y(0)]$. When $X$ is continuous, as is often the case with environmental exposures, there is no longer just one quantity which completely specifies the average effect of interventions on $X$. One might choose to estimate the entire exposure-response curve $E[Y(x)]$ as a function of $x$, but it can be convenient to have a more concise characterization of the effect of an intervention. As we will describe in Section 4, this is typically accomplished in existing approaches by assuming linear and homogeneous treatment effects so that any contrast $E[Y(x)-Y(x^{\prime})]$ is a multiple of the exposure coefficient in a linear regression. However, when the operative parametric assumptions do not hold, this approach does not necessarily result in interpretable causal effects. While estimation of the exposure-response curve using robust causal methods of the kind we advocate is feasible (e.g., Kennedy et al. (2017)), inference is not straightforward, and functional estimation may be unnecessarily complicated if a low-dimensional effect summary is sufficient.
We define a “shift intervention” as a special case of the class of interventions considered in Haneuse and Rotnitzky (2013) and Díaz Muñoz and van der Laan (2012). A shift intervention is an intervention that assigns to individual $i$ with exposure $X_{i}$ the new exposure $X_{i}+\delta$ for some constant $\delta$ (the “shift”). Thus the expected effect of the shift intervention on individual $i$ is $E[Y_{i}(X_{i}+\delta)-Y_{i}(X_{i})]$; notably, this quantity depends on the observed exposure $X_{i}$. If we are interested in population-level effects, we may define $\Delta=E[Y(X+\delta)-Y(X)]$, simply taking the population average of individual effects, where the expectation is taken over the observed data distribution of $X$. The shift intervention effect $\Delta$ is the average effect of increasing the exposure of each individual in the population by $\delta$. This estimand allows the effect of the intervention to depend on the existing distribution of exposure in the population. For example, if the exposure only affects the outcome above a certain threshold, shift interventions might have no effect if current population levels are far below the threshold, but high effects if current population levels are near or above the threshold.
@@
The shift positivity assumption is satisfied if, whenever there is positive probability of exposure level $x$ conditional on confounder value $u$, there is also positive probability of exposure level $x+\delta$ conditional on $U=u$. In the case of spatial confounding, if Assumption 3.8 holds, and Assumption 3.7 is satisfied, then shift positivity also holds for $S$; the logic is identical to that of the case of positivity. Alternatively, we can weaken Assumption 3.6 to a version of Assumption 3.8 that replaces $U$ with $S$. Shift positivity may be a more reasonable assumption than positivity for exposures such as air pollution, where there may be enough variation in the exposure over relatively small regions to satisfy shift positivity, but not enough exposure variation over small regions to cover the entire range of possible exposure values. Considering a shift intervention also requires a weaker version of the ignorability assumption (Haneuse and Rotnitzky, 2013) that only enforces ignorability for the specific exposure values under consideration:
A3.9 (Shift ignorability).
For $(x,u)$ in the support of $(X,U)$, $f(Y(x+\delta)|x,u)=f(Y(x+\delta)|x+\delta,u)$, where $f$ denotes the relevant probability density function.